OpenStudy (anonymous):

hi every1, i have an question that im struggling for some time already and i cant find the answer. A cafeteria have 500 clients a day, it sells the coffe for 2,50 dollars, for each 0,05 increase on the price they lose 10 clients, get the maximun profit possible ?

OpenStudy (amistre64):

its in british units; I can see why you was confused

OpenStudy (amistre64):

its linear

OpenStudy (amistre64):

2.50 + .05n = 500 - 10n is what I see

OpenStudy (amistre64):

well; they aint 'equal' in the regards... just proportional

OpenStudy (amistre64):

500 * 2.5 = 1250 450 * 2.75 = 1237.5

OpenStudy (amistre64):

that doesnt look right.....

OpenStudy (amistre64):

1250.00 1249.50 1248.00 1245.50

OpenStudy (amistre64):

well; max profits appear to be at a price of 2.50 and equal 1250

OpenStudy (amistre64):

I know what I was thinking now lol P(n) = (500-10n)(2.50+.05n)

OpenStudy (anonymous):

You've probably confused him more than he initially was, haha.

OpenStudy (amistre64):

:) you might be right

OpenStudy (anonymous):

the way i tought it was going to to be right was trial and error. just like this 500 x 2,25 and then 490 x 2,30 and keep going

OpenStudy (anonymous):

but its grade 10 math so i guess i need to use some formula like renevue ...forumula

OpenStudy (anonymous):

C = 500 - 10*n V = 2.50 + 0.05*n P = C * V \[P = (500-10n)*(2.50+0.05n) = 1250 - 0.5n ^{2}\] We can observe that to P be the maximum value, we need to put n=0. So, the answer is not increase the value of coffe. Ps.: If is in BRZ, 2.50 is a good price for a good coffee and with a good profit. ;-)